# Real-is-Sim: Bridging the Sim-to-Real Gap with a Dynamic Digital Twin for Real-World Robot Policy Evaluation

### v1

@@ IV-C Demonstration Collection
Figure 5: The success rate of an imitation learning policy as a function of training steps. The plot compares policy evaluation in the offline mode of real-is-sim with the online mode of real-is-sim. The high correlation between the two curves demonstrates that the (significantly faster) offline evaluation mode is a reliable proxy for real-world testing, offering a practical alternative for policy assessment. Error bars are 95% confidence intervals.
IV-D Policy Training
Our policy outputs two components: $q_{d}\in\mathbb{R}^{7}$, the desired joint configuration of the robot, and a scalar progress value $p\in[0,1]$, which indicates the relative completion (progress) of the task. Progress is a heuristic with a ground-truth label that is calculated as the relative time elapsed in the demonstration. It assumes that every demonstration ends at the target position. This was shown to work well as a heuristic in [27]. We use it in our work to evaluate the trained policy without needing to engineer a reward. Our policy is conditioned on the current joint configuration of the simulated robot. In the experiment section, state-based policies are also conditioned on an object’s pose encoded as a position and quaternion (a quaternion is used for simplicity because the pose is stored natively in that form and because it’s double cover property did not prove to be a problem for our task). For image-based policies we use a ResNet to encode an image into a 64 dimensional vector which we condition on. For image-based policies, we found that random crops were crucial for successful learning. Embodied Gaussians allows us to render images from any viewpoint and create virtual cameras in the setup. The representations used to condition the policy—including object pose and rendered images—are explored further in Section V-D. Our policies output 32 actions which are meant to be executed at 30Hz. During evaluation, we execute all 32 actions before resampling.
IV-E Offline Policy Evaluation
During offline policy evaluation, we create 20 parallel environments as shown in Figure 3(b) and run the policies for a predetermined time (90 seconds). We regard any policy that has not reached a progress of above 0.9 as a failure. For state-based policies with 20 environments, we can run the simulation with policy evaluation at 8x realtime on an NVIDIA 6000Ada. With one virtual camera rendering the scene in every environment, the speed is 5x realtime.
V EXPERIMENTS
Figure 6: Success rates of different demonstration collection strategies. A dataset consisting of 30 real demonstrations achieves a success rate of approximately 58%. Since Real-to-Sim can also operate offline, the demonstrations can be augmented with additional simulated data. A combined dataset of 60 demonstrations (30 real and 30 simulated) achieves an 80% success rate. In comparison, a dataset of 60 real demonstrations alone achieves a slightly lower success rate of 72%. Error bars are 95% confidence intervals.
Figure 7: Success rates of policies using different representations to condition the flow matching policy. The baseline image-based policy (not using real-is-sim) uses a static camera and has a success rate of 63%. The state-based policy, which uses the simulator’s object pose achieves a higher success rate of 76%. Lastly, policies trained on the rendered virtual cameras (one statically mounted and another mounted on the gripper) achieve 73% and 82% respectively. This shows that a gripper-mounted camera is likely a better representation for the PushT task. Error bars are 95% confidence intervals.
V-A Experimental Setup
@@
We evaluate our framework on the PushT task: a long horizon task where a robot uses a pusher to position a T-shaped object into a predefined central location. The T-Block’s geometry (represented as a set of spherical shapes attached to a single body frame) is automatically generated using the method from Embodied Gaussians [6]. We use three RealSense D455 cameras running at 90 fps with a resolution of $420\times 270$. The robot mesh is known, and the ground plane is estimated from the camera-generated pointclouds. Gaussians attached to the robot, table, and T-Block are learned once using the Gaussian Splatting procedure outline in [6] and reused across experiments.
@@
For state-based policies, we use the pose of the object encoded as a translation vector $\mathbf{t}\in\mathbb{R}^{3}$ and a unit quaternion $\mathbf{q}\in\mathbb{SO}(3)$, where $\mathbf{t}=(x,y,z)$ and $\mathbf{q}=(q_{x},q_{y},q_{z},q_{w})$. For image-based policies, we use a 64 dimensional feature vector output by a Resnet [22]. The images input to the Resnet are rendered from the simulation using Gaussian Splatting from virtual cameras. These cameras can be placed anywhere. In our experiments, we use a static camera placed in the same location as the real camera and a virtual gripper camera which we mount to the simulated robot’s end-effector (these are shown in Figure 7).
@@
To measure policy success on the PushT task, we define 20 in-distribution starting poses on the table that were not part of the demonstration set. In real-world testing, success is determined by whether the policy drives the T-Block to the target location from these poses. Each pose is tested three times, resulting in 60 measurements per success rate evaluation. Error bars in figures represent 95% confidence intervals.
V-B Offline Evaluation
@@
In this experiment, we demonstrate how real-is-sim’s offline mode facilitates policy training and evaluation. A key challenge in imitation learning lies in selecting optimal checkpoints for deployment. This difficulty arises because the loss used in behavior cloning does not reliably indicate policy performance, and real-world robot evaluations require substantial time and resources. Consequently, researchers often default to using the final training checkpoints despite it being suboptimal.
@@
We demonstrate that real-is-sim’s offline mode enables efficient checkpoint selection through rapid evaluation. For a state-based policy trained on 30 demonstrations, we implement a flow-matching policy trained for 50,000 steps (Figure 5). We then compare two evaluation methodologies: (i) offline rollouts using real-is-sim and (ii) real-world online evaluations. Our offline approach initializes environments with identical test cases to online evaluations. The offline evaluation calculates success rates through the policy-reported progress. While we use this metric to maintain generality, practitioners could implement more sophisticated reward functions for their specific tasks.
@@
We selected four representative checkpoints (10k, 15k, 25k, 50k) for online evaluation (each requires around 1 hour to perform) while evaluating more comprehensively through offline mode. Figure 5 reveals strong correlation between evaluation curves, demonstrating that faster offline evaluations (minutes vs. hours) effectively approximate real-world results. Notably, later checkpoints do not consistently outperform earlier ones, suggesting practitioners could use offline evaluation for either early stopping or optimal checkpoint selection. While this experiment uses identical initial states for both evaluation modes to establish correlation, real applications could benefit from randomized initial states during training-phase evaluations.
V-C Offline Data Collection and Augmentation
@@
In this experiment, we demonstrate how real-is-sim’s offline mode can be used to to collect additional demonstrations to increase task performance (Figure 6). We train a state-based policy using 30 real-world demonstrations collected in online mode - which achieves a success rate of around 57%. We also augment the real-world demonstrations using the simulator only (offline mode). To collect the additional demonstrations, we deploy this policy in offline mode across multiple parallel environments and identify failure states where the policy falls into a local minima. These failure states are then presented to the user to provide an additional 30 demonstrations specifically solving the task under these challenging configurations. By combining the original 30 real-world demonstrations with the 30 new simulated demonstrations, we retrain the policy and observe a significant improvement to 80% success rate.
@@
V-D Representation Flexibility
Another major benefit of real-is-sim’s offline mode is its flexibility in extracting diverse representations from the Embodied Gaussians simulation. For instance, it enables access to privileged information—such as precise object poses—that would be difficult to obtain in the real world. Additionally, virtual cameras can be freely positioned within the system, facilitating visuomotor policy learning from multiple perspectives. In our experiments, we demonstrate that policies leveraging different representations—including state-based object poses, a gripper-mounted virtual camera, and a static virtual camera—all successfully solve the task (Figure 7, 4). Notably, these approaches achieve performance comparable to, and in some cases surpassing, the baseline policy that relies solely on real-world images and robot actions.
@@ V-D Representation Flexibility
VI Limitations
@@
While our framework demonstrates promising results, its applicability is inherently constrained by the capabilities of the underlying physical simulator. Scenarios that deviate significantly from the simulator’s modeling assumptions pose challenges to both model accuracy and corrective mechanisms. To address these constraints, future improvements could focus on enhancing simulator fidelity through adaptive parameter estimation (e.g., real-time optimization of physical properties like friction or mass, as explored in works such as [20] and  [30]). While this work focused on the PushT task to validate core contributions, the methodology is extensible to broader manipulation domains. Enabling applications such as multi-object rearrangement, cloth manipulation, or non-prehensile pushing will require tailoring the simulator’s dynamics model and corrective policies to handle domain-specific interactions.
VII Future Work

### v2

@@ 4 EXPERIMENTS
Figure 7: Success rate of an imitation learning policy versus training steps, comparing evaluation on the real setup and in a virtual environment. The consistency between the two curves shows that the faster virtual-only evaluation is a reliable proxy for real-world testing, providing a practical alternative for policy assessment. Error bars represent 95% confidence intervals.
@@
Experimental Setup We evaluate our framework on the PushT task: a long horizon task where a robot uses a pusher to position a T-shaped object into a predefined location. The T-Block’s geometry (represented as a set of spherical shapes attached to a single body frame) is automatically generated using the method from Embodied Gaussians [1]. We use three RealSense D455 cameras running at 90 fps with a resolution of $420\times 270$. The robot mesh is known, and the ground plane is estimated from the camera-generated pointclouds. Gaussians attached to the robot, table, and T-Block are learned once using the Gaussian Splatting procedure outline in [1] and reused across experiments.
@@
For object state-based policies, we use the pose of the object encoded as a translation vector $\mathbf{p}\in\mathbb{R}^{3}$ and a unit quaternion $\mathbf{R}\in\mathbb{SO}(3)$. For image-based policies, we use a 64 dimensional feature vector output by a ResNet [5] which was jointly trained with the policy. The images input to the ResNet are rendered from the simulation using Gaussian Splatting from virtual cameras. These cameras can be placed anywhere. In our experiments, we use a static camera placed in the same location as the real camera and a virtual gripper camera which we mount to the simulated robot’s end-effector (these are shown in Figure 6).
@@
To evaluate policy success on the PushT task, we define 20 starting poses on the table that are not part of the demonstration set. In real-world testing, success is determined by whether the policy moves the T-Block to the target location from these poses. Each pose is tested three times, resulting in 60 measurements per success rate evaluation. In virtual-only testing, success is determined when the policy outputs a progress value above 0.9, a strategy verified in [8], which serves as a good proxy for success without requiring engineered reward values.
@@
Offline Evaluation In this experiment, we demonstrate how real-is-sim’s offline mode facilitates policy training and evaluation. A key challenge in imitation learning lies in selecting optimal checkpoints for deployment. This difficulty arises because the loss used in behavior cloning does not reliably indicate policy performance, and real-world robot evaluations require substantial time and resources. Consequently, researchers often default to using the final training checkpoints despite it being suboptimal.
@@
We demonstrate that real-is-sim’s offline mode enables efficient checkpoint selection through rapid evaluation. For a state-based policy trained on 30 demonstrations, we implement a flow-matching policy trained for 50,000 steps (Figure 7). We then compare two evaluation methodologies: (i) evaluation on only the virtual world (without the real robot attached and the corrective mechanisms engaged) and (ii) real-world online evaluations. Our virtual-only approach initializes environments with identical test cases to the real-world evaluations. The virtual-only evaluation calculates success rates through the policy-reported progress. While we use this metric to maintain generality, practitioners could implement more sophisticated reward functions for their specific tasks.
@@
We selected four representative checkpoints (10k, 15k, 25k, 50k) for real-world evaluation and performed more comprehensive analysis using faster offline evaluation. Figure 7 shows that, while absolute performance may differ, the relative ordering of checkpoints is preserved - making offline evaluation a practical tool for selecting high-performing policies in minutes rather than hours. Notably, later checkpoints do not consistently outperform earlier ones, suggesting practitioners could use offline evaluation for either early stopping or optimal checkpoint selection. While this experiment uses identical initial states for both evaluation modes to establish correlation, real applications could benefit from randomized initial states during training-phase evaluations.
Figure 8: (a) Success rates of different demonstration collection strategies. (b) Success rates of policies using different representations to condition the flow matching policy. Error bars are 95% confidence intervals.
@@
Offline Data Collection and Augmentation In this experiment, we demonstrate how real-is-sim’s offline mode can be used to to collect additional demonstrations to increase task performance (Figure 8). We train a state-based policy using 30 real-world demonstrations collected in online mode - which achieves a success rate of around 57%. We also augment the real-world demonstrations using the simulator only (offline mode). To collect the additional demonstrations, we deploy this policy in offline mode across multiple parallel environments and identify failure states where the policy falls into a local minima. These failure states are then presented to the user to provide an additional 30 demonstrations specifically solving the task under these challenging configurations. By combining the original 30 real-world demonstrations with the 30 new simulated demonstrations, we retrain the policy and observe a significant improvement to 80% success rate.
@@
Representation Flexibility Another major benefit of real-is-sim’s offline mode is its flexibility in extracting diverse representations from the Embodied Gaussians simulation. For instance, it enables access to privileged information—such as precise object poses—that would be difficult to obtain in the real world. Additionally, virtual cameras can be freely positioned within the system, facilitating visuomotor policy learning from multiple perspectives. In our experiments, we demonstrate that policies leveraging different representations—including state-based object poses, a gripper-mounted virtual camera, and a static virtual camera—all successfully solve the task (Fig. 8b, Fig. 6). Notably, these approaches achieve performance comparable to, and in some cases surpassing, the baseline policy that relies solely on real-world images and robot actions.
@@
5 RELATED WORK
Limitations of Real-World Policy Evaluation: Evaluating robotic policies directly in the real world, while reliable, suffers from being labor-intensive, time-consuming [11, 12], and facing scalability bottlenecks due to manual processes [13]. Consequently, researchers explore simulation for safer, repeatable, and scalable evaluation [14, 15, 16, 17], often focusing on improving simulation fidelity [14, 17, 16, 15]. Our framework differs from this approach by integrating a continuously synchronized simulator throughout the entire policy lifecycle (training, evaluation, deployment). This real-time alignment via sensor data enables direct deployment of simulation-trained policies without further tuning and ensures a unified observation and control space for scalable evaluation.
Addressing the Sim-to-Real Gap in Robotic Manipulation: Simulation offers a scalable alternative to real-world evaluation [11], but the sim-to-real gap remains a key challenge. Existing strategies include domain randomization to enhance robustness [18] and domain adaptation to align representations [19], as well as methods focusing on aligning simulation dynamics [20, 21] or using real-to-sim-to-real approaches [22, 23]. Frameworks like DeepMind’s robot soccer also combine neural rendering with domain randomization [24]. While these methods aim to bridge the sim-to-real gap by making policies more robust or adapting representations, our ”real-is-sim” paradigm eliminates this gap entirely. By continuously aligning the simulator with the real world and treating it as the execution environment, we decouple the policy from sim-real discrepancies. The challenge of domain alignment shifts to the simulator itself, which is addressed through real-time updates.
Leveraging World Models for Robotics: Learned world models enable robots to predict future outcomes for planning and control. These range from 2D predictors [25] to more structured 3D physically grounded models, such as those utilizing Gaussian Splatting [26, 1], which allow for real-time tracking and simulated interaction. Our work builds upon this by exploring the integration of such advanced 3D world models within a behavior cloning framework. This aims to harness the advantages of simulation directly in real-world tasks by training policies on the dynamically updated, real-time digital twin provided by these models.
6 CONCLUSIONS
7 Limitations
@@
While our framework demonstrates promising results, its applicability is fundamentally constrained by the fidelity of the underlying physical simulator. Scenarios that diverge significantly from the simulator’s modeling assumptions present challenges for both forward prediction and corrective mechanisms. Future work could address these limitations by improving simulator accuracy through adaptive parameter estimation—for example, real-time optimization of physical properties such as friction or mass, as explored in [26] and [27].
Another limitation of the current framework is its reliance on visual forces to correct the simulator state. This restricts applicability to scenarios where objects are frequently visible and do not undergo prolonged occlusion or manipulation without observation.
Moreover, for Embodied Gaussians to function effectively, visual corrections must not conflict significantly with the dynamics predicted by the physics simulator. The approach assumes that both the simulator and the visual feedback are aligned, that is, they attempt to move the object in compatible directions. When this assumption breaks down and the corrections diverge, synchronization between the simulation and the real world may not be reliable. However, this limitation can be addressed. The synchronization mechanism can be extended with additional corrective signals beyond visual forces. For example, semantic cues, such as correspondence detection via methods like TAPIR [28] or loss augmentation using segmentation masks as in [29], could provide more robust alignment in challenging settings. The current framework was designed to demonstrate the simplest viable instantiation. Future extensions could incorporate these more sophisticated correction signals to broaden applicability across diverse and visually ambiguous environments.
Another limitation lies in the scene initialization process. In this work, we employed a simple scheme that is sensitive to factors such as the number and placement of cameras, object complexity, and visibility. However, recent advances in learned priors for geometry initialization, such as InstantMesh [30] and InstantSplat [31], offer promising alternatives that could replace or enhance our approach, expanding the set of objects that can be initialized effectively.
Although we focused on the PushT task to validate the core contributions, the proposed methodology is broadly applicable to other manipulation domains. Extending this approach to tasks such as multi-object rearrangement, cloth manipulation, or non-prehensile pushing will require adapting the simulator’s dynamics and corrective policies to accommodate more complex interactions.
ACKNOWLEDGMENT
The QUT authors acknowledge continued support from the Queensland University of Technology (QUT) through the Centre for Robotics. Their work was partially supported by the Australian Government through the Australian Research Council’s Discovery Projects funding scheme (Project DP220102398).
Supplementary Material
Real-is-Sim: Bridging the Sim-to-Real Gap with a Dynamic Digital Twin
Appendix A Follower Configuration vs Leader Configuration
We present real-is-sim as a framework that employs a continuously corrected digital twin of the environment to decouple a behavior cloning policy from the real robot. A central design decision in this framework is to configure the real robot as a follower of the simulated robot. In this setup, when the real robot is connected, it directly mirrors the joint states of the simulated counterpart.
This follower design introduces a distinctive characteristic: the real world consistently lags slightly behind the simulated twin. To our knowledge, this temporal asymmetry is unique to Real-is-Sim and implies that the system always possesses a short-horizon preview of future states. While we find this property conceptually interesting, our primary motivation for adopting the follower configuration is practical, it results in a significantly simpler software architecture and reduces the complexity of sim-to-real integration.
Supplementary Figure 1 illustrates this distinction. In both offline and online modes under the follower configuration (our proposed setup), the system interconnections (blue lines) remain constant. In contrast, the leader configuration—where the real robot dictates state and the simulator reacts—requires dynamic changes to the system architecture. Specifically, three data streams must cross the sim-to-real boundary in the leader case:
  1. 1.
    Control inputs from the policy to the real robot,
  2. 2.
    Control inputs from the real robot to the simulator,
  3. 3.
    Observations from the real robot to the simulator.
By comparison, the follower configuration requires only two such streams. This reduction in cross-boundary communication simplifies the overall system and motivates our recommendation of the follower configuration in Real-is-Sim.
Appendix B Benefits of Collecting Demonstrations with the Real Robot
Figure 2: Tracking failure in two policies, one trained on demonstrations collected with the real robot in the loop, and the other trained entirely in simulation. The left image illustrates a failure case where the simulated physics conflicts with the visual correction required to stay in sync with the real world. Since physics acts as a hard constraint, the correction cannot be applied. Collecting demonstrations in online mode helps avoid such failure modes by encouraging behaviors that maintain alignment between the simulation and reality.
In real-is-sim, we advocate collecting demonstrations with the real robot attached (online mode). During data collection, the user observes the state of the simulator overlaid with the real world, allowing them to continuously assess synchronization quality. This enables the user to adapt their behavior not only to complete the task but also to maintain alignment between the simulation and the real environment. Practically, this means knowing when to move quickly or slowly, and how much clearance to give during fast motions to accommodate slight mismatches between the two domains.
Supplementary Figure 2 illustrates an exaggerated case in which demonstrations are collected entirely in simulation and the resulting policy is deployed on the real robot. Although the task is successfully demonstrated in simulation, it is done with insufficient clearance around objects—causing impacts in the virtual world that do not occur in the real one. The perception system’s correction mechanisms are unable to recover from this divergence, leading to a marked increase in tracking errors during real-world execution.
For this reason, we recommend that at least some demonstrations be collected with the real robot in the loop. This allows both the policy and the demonstrator to implicitly learn behaviors that help keep the simulator and real world in sync.
Appendix C Implementation Details
We use an NVIDIA RTX 6000 Ada as our primary compute device. The robot employed is a Franka Emika with 7 degrees of freedom. The high-impedance controller for the real robot utilizes Ruckig [32] to generate joint trajectories with minimal jerk, in conjunction with the Franka Emika’s built-in joint controller. During demonstration collection, the simulated robot is constrained to act within a 2D plane using Pink [33]. Note that this 2D Cartesian controller is only active during demo collection and only controls the simulated robot. The real robot continues to execute a joint impedance controller based on the simulated robot’s joint positions. Interestingly, the policy learns to operate in 2D, even though the joint controller does not impose such a restriction when the policy is deployed.
Appendix D Alterations to Embodied Gaussians
Our adaptation of Embodied Gaussians [1] (built using [7]) differs from the original framework in two key aspects. First, we omit shape-matching constraints for deformable objects, modeling all entities as rigid bodies composed of spherical collision primitives. This results in a rigid body simulator rather than a particle-based one, improving performance from 30Hz to 60Hz in online mode. Second, we treat the robot as a dynamic entity, simulating mass and inertia for its links and joints, in contrast to the original kinematic formulation, where robot motion was purely kinematic. This change enhances simulation stability and allows the simulated robot to react to the environment. We also found it essential to ensure that the bodies comprising the simulated robot are not affected by gravity. Without this, controlling the simulated robot becomes more difficult, as the controller would need to include a gravity compensation term.
Appendix E Future Work